# Gram-like matrix preserving extensions of noncommutative polynomials to sum of Hermitian Squares

### v1

@@ 3.2. Block Matrix Technique
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{r}\\ a_{1}&1&0&\cdots&0\\ a_{2}&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{r}&0&0&\cdots&1\end{matrix}\end{array}\right)\;.$ (25)
The leading principal minors $M_{l}$ of size $l$, $11\leq l\leq r+$ 1, of the submatrix
$\begin{pmatrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{l-1}\\ a_{1}&1&0&\cdots&0\\ a_{2}&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{l-1}&0&0&\cdots&1\end{pmatrix}$ (26)
of the matrix $\widetilde{G}$, as defined in (25), are given as follows:
$M_{l}=\left\{\begin{array}[]{ll}\sum_{j=1}^{r}a_{i}^{2}+1&\mbox{for $l=1$};\\ \\ \sum_{j=l}^{r}a_{i}^{2}+1&\mbox{for $2\leq l\leq r$};\\ \\ 1&\mbox{for $l=r+1$}.\end{array}\right.$
Therefore the submatrix, as in (26), is positive definite by Theorem 2.4. Moreover, as the matrix $G$, as in (24), is positive semidefinite, so is the matrix $\widetilde{G}$ by Lemma 2.7.
@@
$\begin{split}\widetilde{f}&=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}a_{j}\zeta_{j}+\sum_{j=1}^{r}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}a_{i}^{2}+1\\ &=f+\sum_{j=1}^{r}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}a_{i}^{2}+1.\end{split}$ (29)
It can be noted that $\widetilde{f}$ is obtained by adding $12r+$ 1 many terms with $f$. Moreover, if $\zeta_{i}$ is symmetric for any $i$, then we get the same $\widetilde{f}$ by adding even lesser number of terms. Therefore, first part of the assertion now follows from Theorem 3.5, (28) and (29).
It can be observed that any $\widetilde{G}$ obtained by replacing the $(k+1,k+1)$-th entry of the matrix by any $A>\sum_{j=1}^{r}a_{i}^{2}$ will produce another $\widetilde{f}$ satisfying the first part of the assertion. Hence, we have infinitely many choices for such $\widetilde{f}$. ∎
Remark 3.17.
  1. (1)
    It is clear from the construction of $\widetilde{G}$, as in (25), that even if we start with some other positive semidefinite Gram-like matrix $G^{\prime}$ of the SOHS part of $f$, though we get some different positive semidefinite Gram-like matrix of $\widetilde{f}$ but the polynomial $\widetilde{f}$ will still be the same.
  2. (2)
    The matrix $\widetilde{G}$, as in (25), could also have been constructed as follows:
    $\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}A_{1}&a_{1}&a_{2}&\cdots&a_{r}\\ a_{1}&A_{2}&0&\cdots&0\\ a_{2}&0&A_{3}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{r}&0&0&\cdots&A_{r+1}\end{matrix}\end{array}\right)\;,$
    where $\prod_{i=1}^{r+1}A_{i}\geq\sum_{i=1}^{r}a_{i}^{2}$. In this case, though the newly added block is not positive definite in general but it is positive semidefinite.
The next theorem is towards getting an SOHS extension of $f$ not only with all the expected properties as before but also of arbitrarily high degree.
Theorem 3.18.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be as in Theorem 3.16. Then it is always possible to obtain a SOHS $\tilde{f}\in\mathbb{R}\langle\underline{X}\cup X_{n+1}\rangle$ of any possible even degree $\geq{\rm deg}(f)$, by adding at most $22r+$ 2 many terms with $f$, such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of $f$ as a leading principal submatrix.
Proof.
Let us keep the notations as in Theorem 3.16. We consider the matrix $\widetilde{G}$ given as follows:
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{r}&0\\ a_{1}&1&0&\cdots&0&0\\ a_{2}&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ a_{r}&0&0&\cdots&1&0\\ 0&0&0&\cdots&0&1\end{matrix}\end{array}\right)\;.$ (30)
The leading principal minors $M^{\prime}_{l}$ of size $l$, $21\leq l\leq r+$ 2, of the submatrix
$\begin{pmatrix}\sum_{j=1}^{r}a_{i}^{2}+1&a_{1}&a_{2}&\cdots&a_{l-2}&a_{l-1}\\ a_{1}&1&0&\cdots&0&0\\ a_{2}&0&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ a_{l-2}&0&0&\cdots&1&0\\ a_{l-1}&0&0&\cdots&0&1\end{pmatrix}$ (31)
of the matrix $\widetilde{G}$, as defined in (30), are given as follows:
$M^{\prime}_{l}=\left\{\begin{array}[]{ll}\sum_{j=1}^{r}a_{i}^{2}+1&\mbox{for $l=1$};\\ \\ \sum_{j=l}^{r}a_{i}^{2}+1&\mbox{for $2\leq l\leq r$};\\ \\ 1&\mbox{for $l=r+1$};\\ \\ 1&\mbox{for $l=r+2$ (as $a_{r+1}=0$)}.\end{array}\right.$
Therefore, the submatrix, as in (31), is positive definite by Theorem 2.4. Moreover, as the matrix $G$ is positive semidefinite, so is the matrix $\widetilde{G}$ by Lemma 2.7.
Let $p$ be any positive integer such that $2p\geq{\rm deg}(f)$. We also define $\widetilde{W_{k+r+2}}$ as follows.
$\widetilde{W_{k+r+2}}:=\left(\begin{array}[]{@{}c}\begin{matrix}W_{k}^{t}&1&\zeta_{1}&\cdots&\zeta_{r}&X_{n+1}^{p}\end{matrix}\end{array}\right)^{t}\;.$ (32)
Then the SOHS extension $\widetilde{f}$ of $f$, defined as
$\widetilde{f}:=\widetilde{W_{k+r+2}}^{\ast}\;\widetilde{G}\;\widetilde{W_{k+r+2}},$
where $\widetilde{G}$ and $\widetilde{W_{k+r+2}}$ are as in (30) and (32) respectively, satisfies the assertion. ∎
@@ 3.3. Positive semidefinite completion problem and SOHS via graphs
Now we describe what one means by incompleteness of a symmetric noncommutative polynomial w.r.t its partially specified Gram matrix and its completion.
@@
  1. (1)
    Let $W_{s(d,n)}$ be a column vector of size $s(d,n)$ of all monomials from $\mathbb{R}\langle\underline{X}\rangle$ of degree at most $d$ taken in lexicographic order and $G$ a $s(d,n)\times s(d,n)$ partial positive semidefinite matrix. Then for any positive semidefinite completion $\overline{G}$ of $G$, the polynomial $W_{s(d,n)}^{\ast}\;\overline{G}\;W_{s(d,n)}$ is called a SOHS completion of the polynomial $f:=W_{s(d,n)}^{\ast}\;G\;W_{s(d,n)}$ w.r.t partial positive semidefinite Gram matrix $G$, and is denoted by $\overline{f}^{\;G}$.
  2. (2)
@@
  1. (1)
    Let $f$ be a incomplete polynomial w.r.t a given representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2). Then $f$ is said to be quasi SOHS w.r.t $\mathbf{R}_{f}$ if for any $K\subset\{1,2,\ldots,s(d,n)\}$ such that $b_{ij}$s are known for all $i,j\in K$,
    is a SOHS.
  2. (2)
    Let $f$ be a incomplete polynomial w.r.t a given representation $\mathbf{R}_{f}$, as in (10), along with conditions as in (11). Then $f$ is said to be quasi SOHS w.r.t $\mathbf{R}_{f}$ if for any $K\subset\{1,2,\ldots,k\}$ such that $b_{ij}$s are known for all $i,j\in K$,
@@
  1. (1)
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (1), along with conditions as in (2). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f}}$ w.r.t the partial positive semidefinite Gram matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the specification graph of $G_{f}$ is chordal.
  2. (2)
@@
(1)   Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t a representation
$\mathbf{R}_{f}\;\;:\;\;f=\sum_{i=1}^{s(d,n)}a_{i}\zeta_{i}^{\ast}\zeta_{i}+\sum_{1\leq i<j\leq s(d,n)}b_{ij}(\zeta_{i}^{\ast}\zeta_{j}+\zeta_{j}^{\ast}\zeta_{i}),$
where $a_{i},b_{ij}\in\mathbb{R}$, $\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle$ are all monomials of degree at most $d$. Therefore, for any $K\subset\{1,2,\ldots,s(d,n)\}$ such that $b_{ij}$s are known for all $i,j\in K$,
@@
is positive semidefinite by Theorem 3.5. Therefore any specified principal minor of the Gram matrix $G_{f}$, as in (4), associated with $\mathbf{R}_{f}$ is positive semidefinite, that is to say $G_{f}$ is a partial positive semidefinite matrix. Therefore, it makes sense to talk about SOHS completion of $f$ w.r.t $G_{f}$ following Definition 3.27. Now, the assertion follows from Theorem 3.25.
(2)   Follows similarly. ∎
@@
    Let $\mathcal{G}$ be nonchordal graph. Then also there might exist a quasi SOHS polynomial $f$ with $\mathcal{G}$ as the specification graph of its Gram matrix $G_{f}$ which has a SOHS completion. For example, let $\zeta_{1}=1$, $\zeta_{2}=X_{1}X_{2}$, $\zeta_{3}=X_{2}X_{1}$ and $\zeta_{4}=X_{2}^{2}$ and $W_{4}=\begin{pmatrix}1&X_{1}X_{2}&X_{2}X_{1}&X_{2}^{2}\end{pmatrix}^{t}$. Consider the following polynomial $f$ given as
@@
We now go through a couple of applications of Theorem 3.29.
@@ 3.4. Positive semidefinite completion problem and SOHS via algebraic sets
Unlike commutative setup, the relation between nonnegative noncommutative polynomials and SOHS polynomials is simple. In fact, Helton, in 2002, proved that any positive noncommutative polynomial is a SOHS polynomial (cf. [16]). But the equality between the nonnegative quadratic forms and SOS modulo a suitable projective algebraic set has the potential to allow a SOHS extension of a quasi SOHS polynomial of any even degree. We encode this in the following theorem, which is nothing but a reinterpretation of Theorem 3.29 through the lens of algebraic geometry and commutative algebra. Before that, we quickly recall a couple of definitions along with their notations.
@@
    1. (a)
      Any such $f$ has a SOHS completion $\overline{f}^{\;G_{f}}$ w.r.t the partial positive semidefinite Gram matrix $G_{f}$ associated with $\mathbf{R}_{f}$.
    2. (b)
      The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{s(d,n)-1})$ of $G_{f}$ satisfies $P_{V_{G_{f}}}=\Sigma_{V_{G_{f}}}$.
    3. (c)
      The subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{s(d,n)-1})$ of $G_{f}$ is $2$-regular.
@@
Consider any quasi SOHS polynomial $f$ whose partial positive semidefinite Gram-like matrix $G_{f}$ has the following specification graph $\mathcal{G}_{f}$ :
@@
Then the corresponding subspace arrangement $V_{G_{f}}$ is the projective algebraic subset of $\mathbb{P}^{4}$ whose defining ideal is the homogeneous ideal $I:=\langle x_{0}x_{2},x_{1}x_{2},x_{2}x_{3}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $I$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$-module, as follows :
@@
0   1   2total :3   3   12:3   3   $\begin{split}&0\;\;\;1\;\;\;2\\ \text{total :}\;&3\;\;\;3\;\;\;1\\ 2:\;&3\;\;\;3\;\;\;1\end{split}$
Therefore, we obtain that the ideal $I$ is $2$-regular and so is the projective algebraic set $V_{G_{f}}$. Therefore, by Theorem 3.40, any such quasi SOHS polynomial $f$ admits a SOHS completion w.r.t $G_{f}$. In fact, it also follows from Theorem 3.29 as the specification graph $\mathcal{G}_{f}$ is chordal.
We now apply Theorem 3.29 and Theorem 3.40 to provide a couple of examples of some quasi SOHS polynomials that do not admit SOHS completions.
@@
Example 3.44.
Consider any partial positive semidefinite matrix $G$ which has the following specification graph $\mathcal{G}$ :
[…]
Then the corresponding subspace arrangenment $V_{G}$ is the projective algebraic set sitting inside $\mathbb{P}^{4}$ whose defining homogeneous ideal $J$ is the ideal $\langle x_{0}x_{2},x_{1}x_{3},x_{1}x_{4},x_{2}x_{4}\rangle$. Using Macaulay2 software (cf. [14]), we compute the free resolution and Betti table of $J$, considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$-module, as follows :
Free resolution of $J$ : Denoting $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$ by $R$, we have :
[…]
Betti table of $J$:
0   1   2total :4   4   12:4   3.3:.   1   $\begin{split}&0\;\;\;1\;\;\;2\\ \text{total :}\;&4\;\;\;4\;\;\;1\\ 2:\;&4\;\;\;3\;\;\;\textbf{.}\\ 3:\;&\textbf{.}\;\;\;1\;\;\;1\\ \end{split}$
Therefore, we obtain that the ideal $J$ is $3$-regular and so is the projective algebraic set $V_{G}$. Therefore, by Theorem 3.40, there exists a quasi SOHS polynomial that doesn’t admit a SOHS completion w.r.t $G$. In fact, it also follows from Theorem 3.29 as the specification graph $\mathcal{G}$ contains $\mathcal{C}_{4}$ as induced subgraph.
@@
The flexibility of a Gram-like matrix over a Gram matrix is quite evident from these above examples. As we deal with Gram-like matrices, we could play around with graphs having any number of vertices. For a Gram matrix, we can only deal with graphs with $s(d,n)$ many vertices for some $d\geq 0$ and $n>0$.
@@
As a next step, it would be interesting to generate algorithms out of the criteria provided in Theorem 3.16, Theorem 3.29 and Theorem 3.40 and use that further to produce SDP corresponding to these extension problems.

### v2

@@ 2. Case 2
$\widetilde{G}:=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&\mbox{\Large 0}\\ \hline\cr\mbox{\Large 0}&\begin{matrix}\sum_{j=1}^{r}d_{i}^{2}+1&d_{1}&d_{2}&\cdots&d_{r}\\ d_{1}&1&0&\cdots&0\\ d_{2}&0&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ d_{r}&0&0&\cdots&1\end{matrix}\end{array}\right)\;,$ (13)
where
$d_{i}=\left\{\begin{array}[]{ll}a_{i}&\mbox{if $\zeta_{i}^{\ast}\neq\zeta_{i}$}\;;\\ \\ \tfrac{a_{i}}{2}&\mbox{if $\zeta_{i}^{\ast}=\zeta_{i}$}\;.\end{array}\right.$
The lower diagonal block of the matrix $\widetilde{G}$, as in (13), is positive semidefinite by Lemma 2.11. Moreover, as the matrix $G_{h}$, as in (12), is positive semidefinite, so is the matrix $\widetilde{G}$ by Lemma 2.9.
@@
$\begin{split}\widetilde{f}&=\sum_{i=1}^{m}g_{i}^{\ast}g_{i}+\sum_{j=1}^{r}d_{j}\zeta_{j}+\sum_{j=1}^{r}d_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}d_{i}^{2}+1\\ &=f+\sum_{\zeta_{j}^{\ast}\neq\zeta_{j}}a_{j}\zeta_{j}^{\ast}+\sum_{j=1}^{r}\zeta_{j}^{\ast}\zeta_{j}+\sum_{j=1}^{r}d_{i}^{2}+1.\end{split}$ (16)
It can be noted that $\widetilde{f}$ is obtained by adding at most $12r+$ 1 many terms with $f$. Therefore, first part of the assertion now follows from Theorem 3.5, (15) and (16).
It can be observed that any $\widetilde{G}$ obtained by replacing the $(k+1,k+1)$-th entry of the matrix by any $A\geq\sum_{j=1}^{r}d_{i}^{2}$ will be positive semidefinite by Lemma 2.11 and hence will produce another $\widetilde{f}$ satisfying the first part of the assertion. Hence, we have infinitely many choices for such $\widetilde{f}$. ∎
@@ 3.3. Positive semidefinite completion problem and SOHS via graphs
Now we describe what one means by incompleteness of a symmetric noncommutative polynomial w.r.t its partially specified Gram-like matrix and its completion.
@@
Let $f$ be an incomplete polynomial w.r.t a given representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then $f$ is said to be quasi SOHS w.r.t $\mathbf{R}_{f}$ if for any $K\subset\{1,2,\ldots,k\}$ such that $b_{ij}$s are known for all $i,j\in K$,
@@
Let $f$ be a quasi SOHS polynomial of degree $2d$ w.r.t a representation
where $a_{i},b_{ij}\in\mathbb{R}$, $\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle$ are all monomials of degree at most $d$. Therefore, for any $K\subset\{1,2,\ldots,k\}$ such that $b_{ij}$s are known for all $i,j\in K$,
@@
is positive semidefinite by Theorem 3.5. Therefore any specified principal minor of the Gram-like matrix $G_{f}$, associated with $\mathbf{R}_{f}$ is positive semidefinite, that is to say $G_{f}$ is a partial positive semidefinite matrix. Therefore, it makes sense to talk about SOHS completion of $f$ w.r.t $G_{f}$ following Definition 3.33. Now, the assertion follows from Theorem 3.31. ∎
@@
  1. (1)
    It is easy to observe that Theorem 3.35 holds for Gram matrix as well, by taking $k=s(d,n)$. To do so, one only has to analogously define Definition 3.32, Definition 3.33 and Definition 3.34 w.r.t a representation as $\mathbf{R}_{f}$, as in (1), along with conditions in (2).
@@
    Let $\mathcal{G}$ be a nonchordal graph. Then also there might exist a quasi SOHS polynomial $f$ with $\mathcal{G}$ as the specification graph of its Gram matrix $G_{f}$ which has a SOHS completion. For example, let $\zeta_{1}=1$, $\zeta_{2}=X_{1}X_{2}$, $\zeta_{3}=X_{2}X_{1}$ and $\zeta_{4}=X_{2}^{2}$ and $W_{4}=\begin{pmatrix}1&X_{1}X_{2}&X_{2}X_{1}&X_{2}^{2}\end{pmatrix}^{t}$. Consider the following polynomial $f$ given as
@@
We now go through some applications of Theorem 3.35.
@@
Corollary 3.40.
as in (8) along with conditions as in (9). If out of $\tfrac{k(k-1)}{2}$ many $b_{ij}$s, at most $3$ are specified, then $f$ has a SOHS completion.
Proof.
The hypothesis ensures that the specification graph $\mathcal{G}_{G_{f}}$ of the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ has $k$ vertices and at most $3$ edges. Therefore, $\mathcal{G}_{G_{f}}$ can’t have $\mathcal{C}_{n}$ as an induced subgraph for $n\geq 4$ and $\mathcal{G}_{G_{f}}$ is chordal. Therefore it follows from Theorem 3.35. ∎
@@ 3.4. Positive semidefinite completion problem and SOHS via algebraic sets
Unlike commutative setup, the relation between positive noncommutative polynomials and SOHS polynomials is simple. In fact, any positive noncommutative polynomial is a SOHS polynomial (cf. Remark 2.3). But the equality between the nonnegative quadratic forms and SOS modulo a suitable projective algebraic set has the potential to allow a SOHS extension of a quasi SOHS polynomial of any even degree. We encode this in the following theorem, which is nothing but a reinterpretation of Theorem 3.35 through the lens of algebraic geometry and commutative algebra. Before that, we quickly recall a couple of definitions along with their notations.
@@
Consider any quasi SOHS polynomial $f$ whose partial positive semidefinite Gram-like matrix $G_{f}$ has the following disconnected specification graph $\mathcal{G}_{f}$ :
@@
Then the corresponding subspace arrangement $V_{G_{f}}$ is the projective algebraic subset of $\mathbb{P}^{4}$ whose defining ideal is the homogeneous ideal $I:=\langle x_{0}x_{2},x_{0}x_{3},x_{0}x_{4},x_{1}x_{2},x_{1}x_{3},x_{1}x_{4},x_{2}x_{4},x_{3}x_{4}\rangle$. Using Macaulay2 software, we compute the free resolution and Betti table of $I$ considered as a $\mathbb{R}[x_{0},x_{1},x_{2},x_{3},x_{4}]$-module, as follows :
@@
0    1    2   3total :8   14   9   22:8   14   9   $\begin{split}&0\;\;\;\;1\;\;\;\;2\;\;\;3\\ \text{total :}\;&8\;\;\;14\;\;\;9\;\;\;2\\ 2:\;&8\;\;\;14\;\;\;9\;\;\;2\\ \end{split}$
Therefore, we obtain that the ideal $I$ is $2$-regular and so is the projective algebraic set $V_{G_{f}}$. Therefore, by Theorem 3.47, any such quasi SOHS polynomial admits a completion w.r.t $G_{f}$. In fact, it also follows from Theorem 3.35 as the specification graph $\mathcal{G}_{f}$ is chordal.
We now apply Theorem 3.35 and Theorem 3.47 to provide an example of some quasi SOHS polynomials that do not admit SOHS completions.
@@
The flexibility of a Gram-like matrix over a Gram matrix is quite evident from these examples. As we deal with Gram-like matrices, we could play around with graphs having any number of vertices. For a Gram matrix, we can only deal with graphs with $s(d,n)$ many vertices for some $d\geq 0$ and $n>0$.
@@
As a next step, it would be interesting to generate algorithms out of the criteria provided in Theorem 3.19, Theorem 3.22, Theorem 3.24, Theorem 3.35 and Theorem 3.47 and use that further to produce SDP corresponding to these extension problems.